what is the soultion to this simultaneous equation using substion or addtion method
x+2y=15
x-2y=-9

Answer by sarah_adam(201)

x+2y=15 --- eq 1
x-2y=-9 --- eq 2
-------
2x = -6 (adding eq1 and eq2)
x = -6/2 = -3
substitute value of x in eq 1
2y = 15 - x
2y = 15 - (-3)
2y = 15+3 = 18
y = 18/2 = 9
x = -3; y = 9